# LAPD: Langevin-Assisted Bayesian Active Learning for Physical Discovery

### v1

@@ 5.2 Burgers’ Equation with Active Learning
We next evaluate LAPD on Burgers’ Equation. The original data here distributed around 4001 positions equally spaced in [-8,8] and 1001 different time steps equally spaced in [0,10]. The initial condition is set to the same as in Section 4.3. Among the original data, 3 different time steps are selected as the training time step. Here, we chose 1s,5s, and 8s, and actually, this time step can be chosen randomly, and even the whole training dataset can be chosen randomly throughout the whole dataset. Under this scenario, the calculation of partial derivatives with respect to $t$ can be difficult to obtain, hence in need of active learning. Again, in this synthetic data scenario, we calculate $\mathbf{U}_{t}$ based on the original data. The candidate library adopted in this experiment is the same as in Section 4.3 with 10 terms $[1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. Here we inject i.i.d. noise on $\dot{\mathbf{U}}$ following $lognormal(0,0.1)$.
@@
Training begins with an initial set of 20 randomly selected data. In each subsequent round, 10 additional data are selected via the specified acquisition functions and added to $\mathcal{M}$. After the first round of training with the initial data, the minimum $d(\mathbf{u})$ is $6.769\times 10^{-11}$, while the maximum $d(\mathbf{u})$ is $1.443$, indicating a significant relative difference of more than $10^{10}$, so density adjustment is performed. Here, we adopt $\lambda=0.5,\alpha=0.3$. As shown in Figure 8, AL with a hybrid acquisition function performs better, requiring fewer training points. The threshold Tol is set to 0.3, and experiments with other acquisition functions are carried out for the same number of rounds as the uncertainty+space-filling acquisition function. The hybrid method identifies the correct candidate functions with 50 points, while the Uncertainty method and the Random method are left with only 4 and 7 basis functions, respectively. For the error bar to drop below $10^{-2}$, the random sampling strategy requires 90 data, whereas the hybrid strategy only requires 50 data, reducing the required measurements by 44%. As shown in Table 6, the random sampling approach results in an equation contaminated with extraneous terms such as $u,u^{2},uu_{x},u_{x}u_{xx}$, which deviate significantly from the true system. This suggests that randomly selected training points fail to capture the underlying physics, leading to overfitting or the inclusion of spurious terms. On the other hand, uncertainty-based sampling shows improvement by reducing the number of incorrect terms. However, the coefficient estimates remain suboptimal, particularly for the nonlinear interaction term $u\frac{\partial u}{\partial x}$. This indicates that an uncertainty-only acquisition function may still suffer from biased selection, potentially over-exploring regions with high variance while neglecting global representativeness. The hybrid uncertainty+space-filling strategy demonstrates the most accurate identification of the governing equation. The identified terms match the true system structure, and the coefficient values closely approximate their theoretical counterparts. This highlights the advantage of balancing exploration (uncertainty) and diversity (space-filling) when selecting training points, leading to more robust equation discovery with fewer data points.
@@
Figure 8: Active Learning on Burgers’ Equation: Log-scale plot of the error bar for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added.
@@
Table 6: Comparison of true system (Burgers’ Equation) and identified systems for different strategies with 70 training data
System $\frac{\partial u}{\partial t}$
True system $\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-0.1\frac{\partial^{2}u}{\partial x^{2}}=0$
Random $\frac{\partial u}{\partial t}-0.054u-0.075\frac{\partial u}{\partial x}+0.147\frac{\partial^{2}u}{\partial x^{2}}+0.944u^{2}+0.060u\frac{\partial u}{\partial x}+0.115\frac{\partial u}{\partial x}\frac{\partial^{2}u}{\partial x^{2}}=0$
Uncertainty $\frac{\partial u}{\partial t}+0.14\frac{\partial u}{\partial x}+0.714u\frac{\partial u}{\partial x}-0.073\frac{\partial^{2}u}{\partial x^{2}}=0$
@@
Uncertainty+Space-Filling $\frac{\partial u}{\partial t}+0.971u\frac{\partial u}{\partial x}-0.090\frac{\partial^{2}u}{\partial x^{2}}=0$
@@ 6 Conclusion and Discussion
This work presents a novel Lagevin-assisted Bayesian active learning framework, LAPD, which integrates reSGLD with AL to discover governing equations from experimental data. Our comprehensive evaluation across diverse dynamical systems demonstrates that incorporating reSGLD sampling within the physical discovery process provides significant advantages over traditional approaches, especially in the face of noise. By simultaneously operating multiple Markov chains at different temperatures, LAPD achieves a faster exploration of the parameter space, effectively navigating the complex landscapes of nonlinear dynamical systems. Traditional sparse regression methods for physical discovery (like SINDy) provide point estimates of system coefficients but offer limited insight into model confidence intervals. In contrast, LAPD offers robust UQ for both model parameters and system trajectories. Moreover, while sparsity-inducing priors (as in UQ-SINDy) can theoretically promote sparse solutions, our experiments improved robustness and accuracy when identifying system dynamics and estimating the uncertainties. Furthermore, our AL extension introduces a hybrid acquisition function that balances uncertainty reduction with domain coverage. This approach significantly outperforms both random sampling and pure uncertainty-based acquisition, reducing measurement requirements by 44-57% across test cases.
While the LAPD demonstrates robust performance across our test cases, several methodological challenges remain. First, the computational cost of reSGLD sampling exceeds that of deterministic optimization approaches. For very large systems or real-time applications, this overhead may become prohibitive. Algorithmic optimizations like adaptive temperature scheduling or parallel chain implementation could address this limitation. Second, our current implementation determines the initial library of candidate functions a priori. For complex systems where appropriate basis functions are unknown, this presents a practical challenge. Future work could integrate automatic basis function discovery, perhaps by incorporating neural networks to learn transformed coordinates where dynamics admit sparse representations.
In summary, LAPD represents a significant step toward robust, uncertainty-aware physical discovery from limited data. By bridging Bayesian inference, stochastic sampling, and AL, it addresses fundamental challenges in data-driven science. As experimental data in many fields remain costly and imperfect, methods like LAPD that maximize information extraction while quantifying uncertainty will become increasingly valuable for scientific discovery.

### v2

@@ 5.2 Burgers’ Equation with AL
We next evaluate BLADE with AL on Burgers’ equation. The original data here distributed approximately 4,000 positions equally spaced in [-8,8] and 1,000 different time steps equally spaced in [0,10]. The initial condition is set to the same as in Section 4.3. The pool of potential design points is constructed by selecting three distinct time steps (1s, 5s, and 8s) from the original data. These time points serve as a representative random sample, and any arbitrary selection of time steps from the dataset would be equally suitable. Under this scenario, the calculation of partial derivatives with respect to $t$ can be difficult to obtain, hence in need of active learning. Again, in this synthetic data scenario, we calculate $\mathbf{U}_{t}$ based on the original data. The candidate library adopted in this experiment is the same as in Section 4.3 with 10 terms $1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. Here we inject i.i.d. noise on $\dot{\mathbf{U}}$ following lognormal(0,0.1)lognormal\,(0,0.1).
@@
Training begins with an initial set of 20 randomly selected data. In each subsequent round, 10 additional data are selected via the specified acquisition functions and added to $\mathcal{M}$. The tolerance of convergence Tol is set to 0.3, and experiments with other acquisition functions are carried out for the same number of rounds as the uncertainty+space-filling acquisition function. Following the initial training phase, the values of $d(\mathbf{u})$ ranged from a minimum of $6.769\times 10^{-11}$ to a maximum of $1.443$. This substantial relative difference, exceeding ten orders of magnitude, prompted a density adjustment with parameters set to $\lambda=0.5,\alpha=0.3$. As shown in Figure 8, AL approach utilizing a hybrid acquisition function demonstrates enhanced data efficiency, achieving strong performance with a reduced number of training points. The hybrid method identifies the correct candidate functions with 50 points, while the uncertainty method and the random method are left with 4 and 7 basis functions, respectively. For the Total Error Bar to drop below $10^{-2}$, the random sampling strategy requires 90 data, whereas the hybrid strategy only requires 50 data, which reduces the required measurements by 44%. As shown in Table 6, the random sampling approach results in an equation contaminated with extraneous terms such as $u,u^{2},uu_{x},u_{x}u_{xx}$, which deviate significantly from the true system. This suggests that randomly selected training points may not be able to adequately capture the underlying physics, leading to overfitting or the inclusion of spurious terms. On the other hand, uncertainty-based sampling shows improvement by reducing the number of incorrect terms. However, the coefficient estimates remain suboptimal, particularly for the non-linear interaction term $u\frac{\partial u}{\partial x}$. This indicates that an uncertainty-only acquisition function may still suffer from biased selection, potentially over-exploring regions with high variance while neglecting global representativeness. The hybrid uncertainty+space-filling strategy demonstrates the most accurate identification of the governing equations. The identified terms match the true system structure, and the coefficient values closely approximate their theoretical counterparts. This highlights the advantage of balancing exploration (uncertainty) and diversity (space-filling) when selecting training points, leading to more robust equation discovery with fewer data points.
@@
Figure 8: Active Learning on Burgers’ equation: Log-scale plot of the Total Error Bar for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added.
@@
Table 6: Comparison of true system (Burgers’ equation) and identified systems for different strategies with 70 training data. Our method (Uncertainty+Space-Filling) is shown bolded.
@@
Methods Learned systems True system $\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-0.1\frac{\partial^{2}u}{\partial x^{2}}=0$ Random $\frac{\partial u}{\partial t}-0.054u-0.075\frac{\partial u}{\partial x}+0.147\frac{\partial^{2}u}{\partial x^{2}}+0.944u^{2}+0.060u\frac{\partial u}{\partial x}+0.115\frac{\partial u}{\partial x}\frac{\partial^{2}u}{\partial x^{2}}=0$ Uncertainty $\frac{\partial u}{\partial t}+0.14\frac{\partial u}{\partial x}+0.714u\frac{\partial u}{\partial x}-0.073\frac{\partial^{2}u}{\partial x^{2}}=0$ Uncertainty+Space-Filling $\frac{\partial u}{\partial t}+0.971u\frac{\partial u}{\partial x}-0.090\frac{\partial^{2}u}{\partial x^{2}}=0$
5.3 Convection-Diffusion Equation with AL
We further evaluate BLADE with active learning on the convection–diffusion equation. We randomly select 600 points from the full dataset in Section 4.4 to form the pool of potential design points, denoted by $\mathcal{D}$. Because these points are sampled irregularly in $(x,t)$, accurate estimation of partial derivatives with respect to $x$ and $t$ becomes more challenging, which highlights the need for active learning. As in the previous synthetic-data experiments, we compute $\mathbf{U}_{t}$ using the original data from Section 4.4. The candidate library is the same as in Section 4.4, consisting of 10 terms: $1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. We additionally perturb $\mathbf{U}$ with i.i.d. additive Gaussian noise at a level of $0.1\%$.
Training starts from an initial set of 20 randomly selected data points. At each subsequent active learning round, 10 additional points are selected by the specified acquisition function and added to $\mathcal{M}$. The tolerance of convergence is fixed at $\mathit{Tol}=0.3$. To ensure a fair comparison, all acquisition strategies are run for the same number of rounds as the uncertainty+space-filling strategy. For the hybrid acquisition function, we use the density-adjustment parameters $\lambda=0.5$ and $\alpha=0.3$. Figure 9 shows that the hybrid acquisition strategy (uncertainty+space-filling) is substantially more data-efficient than either random sampling or uncertainty-only sampling. In Fig. 9(a), we report the coefficient error, measured as the mean squared error (MSE) between the identified PDE coefficients and the ground-truth coefficients. A slight increase in coefficient MSE is observed from 20 to 30 training points for all three strategies. This behavior is expected in sequential active discovery: early queried points are chosen to improve informativeness and coverage, which can temporarily shift the posterior and the selected sparse support before the coefficient estimates stabilize. As more points are added, the coefficient MSE decreases for all methods, but the reduction is much sharper for the hybrid strategy, dropping by 5 orders of magnitude and reaching the lowest error among all methods. Figure 9(b) reports the total error bar. The hybrid strategy shows the fastest reduction and achieves a near-converged value by 40 training points, while the random and uncertainty-only strategies remain substantially larger at the same budget. This behavior is consistent with the identified systems in Table 7: with 50 training points, the hybrid strategy recovers the correct convection–diffusion structure with coefficients very close to the ground truth ($1.002$ for $u_{x}$ and 0.996-0.996 for $u_{xx}$), whereas the random strategy misses the diffusion term $u_{xx}$ entirely, and the uncertainty-only strategy identifies the correct basis but with noticeably larger coefficient error. Overall, these results highlight the benefit of combining exploration (uncertainty) with diversity (space-filling): the hybrid acquisition function improves support recovery and coefficient accuracy using fewer training points, leading to more robust equation discovery in the irregularly sampled setting.
Figure 9: Illustration of active learning for the BLADE method on the Convection-Diffusion equation. (a) Log-scale plot of the MSE of coefficients for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added. (b) Log-scale plot of the Total Error Bar for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added.
Table 7: Comparison of true system (Convection-Diffusion equation) and identified systems for different strategies with 50 training data. Our method (Uncertainty+Space-Filling) is shown bolded.
Methods Learned systems True system $\frac{\partial u}{\partial t}+\frac{\partial u}{\partial x}-\frac{\partial^{2}u}{\partial x^{2}}=0$ Random $\frac{\partial u}{\partial t}+0.869\frac{\partial u}{\partial x}=0$ Uncertainty $\frac{\partial u}{\partial t}+0.767\frac{\partial u}{\partial x}-0.662\frac{\partial^{2}u}{\partial x^{2}}=0$ Uncertainty+Space-Filling $\frac{\partial u}{\partial t}+1.002\frac{\partial u}{\partial x}-0.996\frac{\partial^{2}u}{\partial x^{2}}=0$
@@ 6 Conclusion and Discussion
This work presents a novel Lagevin-assisted Bayesian active learning framework, BLADE, which integrates reSGLD with AL to discover governing equations from noisy and limited experimental data. Our comprehensive evaluation across diverse dynamical systems demonstrates that incorporating reSGLD sampling within the physical discovery process provides significant advantages over traditional approaches, especially in the face of noise. By simultaneously operating multiple Markov chains at different temperatures, BLADE achieves a better exploration of the parameter space, effectively navigating the complex landscapes of nonlinear dynamical systems. Traditional sparse regression methods for physical discovery (like SINDy) provide point estimates of system coefficients but offer limited insight into model UQ. In contrast, BLADE offers robust UQ for both model parameters and system trajectories. Moreover, while sparsity-inducing priors (as in UQ-SINDy) can theoretically promote sparse solutions, our experiments improved robustness and accuracy when identifying system dynamics and estimating the uncertainties. Furthermore, our AL extension introduces a hybrid acquisition function that balances uncertainty reduction with domain coverage. This approach significantly outperforms both random sampling and pure uncertainty-based acquisition, reducing measurement requirements by 44-57% across test cases.
While BLADE demonstrates strong performance across benchmarks, several challenges remain. First, the computational cost of reSGLD sampling exceeds that of deterministic optimization approaches. For very large systems or real-time applications, this overhead may become prohibitive. Future work would explore algorithmic optimizations, like adaptive temperature scheduling or parallel chain implementation. Second, our current implementation determines the initial library of candidate functions a priori. For complex systems where this library is incomplete or unknown, incorporating neural operators or manifold-learned feature spaces may enable automatic basis discovery. Third, BLADE presently relies on finite-difference derivative estimates, which can be fragile in noisy PDE settings. Pointwise numerical differentiation is well known to amplify measurement noise, which can degrade the candidate library construction and the downstream system-identification step. Improving robustness in high-noise PDE settings is an important direction for future work. One promising approach is to extend BLADE to weak-form or integral formulations, which avoid explicit pointwise derivative estimation by transferring derivatives from noisy data to smooth test functions and thus naturally reduce noise amplification [messenger2021weak, messenger2020weak]. Another effective strategy is transforming the candidate library from the time/spatial domain into the Laplace or frequency domain. Methods such as LES-SINDy [zheng2024les_sindy] employ the Laplace transform alongside integration by parts to evaluate derivatives analytically in the transformed space, effectively handling noisy measurements and high-order derivatives without direct numerical differentiation. Similarly, filtering techniques based on the Discrete Fourier Transform (DFT) can be applied to isolate and attenuate noise-dominated frequencies prior to candidate library construction [thanasutives2022noiseaware]. A third direction is to incorporate neural-network-based denoising or surrogate representations and compute derivatives via automatic differentiation, potentially combined with physics-informed or time-stepping-constrained architectures to better separate measurement noise from the underlying dynamics [rudy2019deep, thanasutives2022noiseaware]. These extensions would decouple BLADE’s Bayesian inference and uncertainty-quantification components from the current finite-difference preprocessing bottleneck, and could substantially improve robustness in strongly noisy regimes.
In summary, BLADE constitutes a robust, uncertainty-aware, and data-efficient framework for the discovery of interpretable dynamical systems. By bridging Bayesian inference, stochastic sampling, and AL, it addresses fundamental challenges in data-driven science. As high-fidelity experimental data remain costly and limited, the principles of Bayesian active discovery embodied in BLADE are expected to play an increasingly central role in the next generation of data-driven physics and engineering.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
Data will be made available on request from the authors.
Funding
We would like to thank the support of National Science Foundation (DMS-2533878, DMS-2053746, DMS-2134209, ECCS-2328241, CBET-2347401 and OAC-2311848), and U.S. Department of Energy (DOE) Office of Science Advanced Scientific Computing Research program DE-SC0023161, the SciDAC LEADS Institute, and DOE–Fusion Energy Science, under grant number: DE-SC0024583.
Appendix A Replica-exchange hyperparameter specification and swap-rate behavior
In the reSGLD sampler used in BLADE, we fix the two temperatures in Eq. (8) as $\tau_{1}=1,\tau_{2}=2$ and treat the ratio $\tilde{\sigma}^{2}/C$ in the corrected swapping function Eq. (9) as an effective hyperparameter. Following prior reSGLD implementations [deng2020non, zheng2024constrained], we tune the ratio $\tilde{\sigma}^{2}/C$ directly, rather than estimating $\tilde{\sigma}$ and choosing $C$ separately. In the experiments, we first conduct a swap-rate ablation on the Burgers equation, where we plot the swap acceptance rate over algorithmic iterations for representative values of $\tilde{\sigma}^{2}/C$; the dashed vertical line in Fig. 10 marks the warm-up boundary at 5000 iterations. This experiment is used to illustrate how the choice of $\tilde{\sigma}^{2}/C$ affects exchange efficiency in practice, and values around $10^{-11}$ lead to stable post–warm-up swap rates in a practically usable range for Burgers. We follow the same tuning principle in the other benchmarks, with $\tilde{\sigma}^{2}/C$ set to $6$ for the Lotka–Volterra system, $3$ for the Lorenz system, and $10^{-11}$ for both the Burgers and convection–diffusion equations, while fixing the replica temperatures to $\tau_{1}=1$ and $\tau_{2}=2$ throughout.
Tuning procedure.
The choice of $\tilde{\sigma}^{2}/C$ is made empirically based on the observed swap acceptance behavior. Practically useful settings typically yield post–warm-up swap acceptance rates in a moderate range of approximately $5\%$ to $15\%$. In practice, we begin with a coarse scan over a wide range of candidate values, typically starting from $10$ and decreasing by orders of magnitude, in order to identify a regime in which replica exchange is neither degenerate nor overly suppressed. We then make local adjustments within that regime when needed for a given benchmark. As a result, selecting $\tilde{\sigma}^{2}/C$ involves a modest amount of empirical tuning in practice.
Effect of $\tilde{\sigma}^{2}/C$ on replica-exchange dynamics.
Figure 10 shows the swap acceptance rate over algorithmic iterations for representative values of $\tilde{\sigma}^{2}/C$ on the Burgers benchmark during the first round of coefficient selection, with the dashed vertical line marking the warm-up boundary at 5000 iterations. The main observation is that replica-exchange behavior remains relatively stable for values near $10^{-11}$, where the post–warm-up swap acceptance rate stays in a moderate range. This suggests that the practical tuning of $\tilde{\sigma}^{2}/C$ is not overly sensitive within a local neighborhood of the selected value. By contrast, substantially larger values suppress exchange more strongly and lead to noticeably lower post–warm-up swap rates. For the Burgers’ equation benchmark, values of $\tilde{\sigma}^{2}/C$ above $10^{-8}$ lead to relatively inaccurate estimation of the coefficient of $u_{xx}$, causing this term to be omitted and thereby resulting in incorrect system discovery. Overall, this behavior supports our empirical choice of $\tilde{\sigma}^{2}/C$ and indicates that the selected value lies in a stable operating range for replica exchange.
Figure 10: Swap acceptance rate over algorithmic iterations for representative values of $\tilde{\sigma}^{2}/C$ on the Burgers’ benchmark (first round). The dashed vertical line indicates the warm-up boundary at 5000 iterations. Values of $\tilde{\sigma}^{2}/C$ around $10^{-11}$ produce stable post–warm-up swap rates in a practically usable range, whereas overly large values substantially suppress exchange.